# Never Miss A Beat: An Efficient Recipe for Context Window Extension of Large Language Models with Consistent “Middle” Enhancement

### v1

@@ Abstract
Recently, many methods have been developed to extend the context length of pre-trained large language models (LLMs), but they often require fine-tuning at the target length ($\gg 4K$) and struggle to effectively utilize information from the middle part of the context. To address these issues, we propose Continuity-Relativity indExing with gAussian Middle (CREAM), which interpolates positional encodings by manipulating position indices. Apart from being simple, CREAM is training-efficient: it only requires fine-tuning at the pre-trained context window (e.g., Llama 2-4K) and can extend LLMs to a much longer target context length (e.g., 256K). To ensure that the model focuses more on the information in the middle, we introduce a truncated Gaussian to encourage sampling from the middle part of the context during fine-tuning, thus alleviating the “Lost-in-the-Middle” problem faced by long-context LLMs. Experimental results show that CREAM successfully extends LLMs to the target length for both Base and Chat versions of Llama2-7B with “Never Miss A Beat”. Our code will be publicly available soon.
@@
Figure 1: Results of applying different position interpolation methods to the “Lost-in-the-Middle” task on CREAM and PoSE [1]. We can see that CREAM outperforms PoSE [1] at every position, with a particularly improvement in the middle.
@@ 1 Introduction
Transformer-based Large Language Models (LLMs) are typically pre-trained with a fixed context window size, e.g., 4K tokens in Touvron et al. [2]. However, many downstream applications, including in-context learning [3, 4] and LLM agents [5, 6]. necessitate the processing of significantly longer contexts, e.g., up to 256K tokens. Recent works have proposed promising approaches that efficiently extend the context window of pre-trained LLMs by interpolating Positional Encodings (PEs) [7, 8, 9, 10, 11] with a short period of fine-tuning. Unlike other techniques such as efficient transformer [12, 13] and memory augmentation [14], PE-based methods do not necessitate alterations to the model’s architecture or the incorporation of supplementary modules. Consequently, PE-based methods offer the advantages of straightforward implementation and rapid adaptation, making them a practical solution for extending the operational range of LLMs in tasks involving larger context windows.
Despite the simplicity and effectiveness, existing PE-based methods exhibit two significant limitations. First, prior approaches, such as positional interpolation [7], still require fine-tuning on the target context window size, which imposes a substantial computational overhead [1]. Secondly, though some PE methods have demonstrated potential in handling extremely long sequences, as evidenced by low sliding window perplexity scores, their performance deteriorates notably in “in-the-middle” scenarios [15]. Specifically, when the model is required to accurately retrieve and process content located in the middle of an extended context, there is a marked drop in performance on the extended window size (Figure 1 and Figure 3).
@@
    CREAM can not only fine-tune within the pre-training context window size, but also alleviate the issue of the model easily getting lost in the middle. e.g., CREAM-YaRN outperforms PoSE-YaRN  [1] by over 20% on average in the “Lost in the Middle” [15] task.
@@
    CREAM can further be enhanced by integrating novel designs on positional interpolation frequencies (such as Linear [7], NTK [8], Yarn [9], etc.), and can be extended to context window sizes of up to 256K or beyond.
@@
    CREAM-Chat model requires only 100 steps of instruction-tuning to achieve nearly perfect performance on the Needle-in-a-Haystack pressure test, and it outperforms existing strong baselines on LongBench [16].
@@ 2.1 Preliminaries
Transformer-based language models typically encode positional indices sequentially as $\{0,1,\cdots,N-1\}$. Traditional length extension methods [7, 8, 9] directly fine-tune on the target length $L$ with an updated positional index. This approach preserves the continuity of all absolute positions and learns all position indices within $[0,L-1]$, thereby successfully extending to the target length. Furthermore, PoSE [1] attributed their superior performance over RandPos [17] to the ensured continuity of segments during fine-tuning.
@@
Relative positional encoding (RPE) [18] has been proposed as an effective positional encoding method, where only the relative positions between two tokens are considered. Similar as prior works [17, 1, 19], our work focuses on rotary positional encoding (RoPE) [20], which is one of the most prominent RPE methods and has been widely applied to LLMs including the recent Llama family [21, 2, 22]. In RoPE, only the relative distances between position pairs $(|j-i|;0\leq i<j\leq L-1)$ are learned during fine-tuning (Appendix A). Due to this property, we can manipulate the position indices such that all relative positions between $[0,L-1]$ are learnable within the pre-trained window size.
@@ 2.2 Proposed Recipe: Continuity-Relativity indExing with gAussian Middle (CREAM)
We first discuss the motivations behind our design of the context length. First, prior works [23, 24] observed that a significant amount of attention score is allocated to the beginning tokens of a sequence, which can potentially encode absolute positional information even without explicit positional encoding [25]. Secondly, the starting and ending tokens of long contexts can be treated as two pointers that localize the middle indices with the help of relative encodings. Therefore, we divide the pre-trained context window into three segments. The detailed ablation results are shown in Section 3.5.
@@
For the sake of continuity, we set the $L_{h}$ and $L_{t}$ to a very small value $k$, where $0\textless k\ll N$. Specifcally, we use $k=32$ in our experiments. This choice allows the middle segment to closely approximate the pre-trained context window. To maintain relativity, we divide $N$ equally into three parts and fix the $L_{h}$ and $L_{t}$ to $N/3$, enabling the model to learn as many relative positions as possible. In our fine-tuning process, both types of examples are sampled with equal probability to maintain balance.
@@
To better focus the training process on the middle part of the long context, we introduce a truncated Gaussian function. This approach reduces the interval overlap in Equation 2 and directs the model’s attention towards the middle section of the long context. In Appendix B, we provide theoretical justifications of our truncated Gaussian design, indicating that the maximization of $|D_{r}|$ holds for middle positions in $[N,L/2)\cup(L/2,L-N]$.
@@ 3.1 Experimental Setup
We use Llama-2-7B and Llama-2-7B-Chat [2] as the base models and extend their pre-trained context window size of 4K to a target context length of 32K. The extended models are referred to as CREAM-Base and CREAM-Chat, respectively. Note that, though the target context length is 32K, we do not have to fine-tune CREAM on 32K token long text (see Section 2.2).
@@
We conduct long-context LLM evaluation of CREAM-Base on LongChat-Lines [26] and Lost-in-the-Middle [15]. Ideally, fine-tuning should not disrupt what the base model has learned, so we further evaluate CREAM-Base on the language modeling task and the evaluation benchmark [27] adopted by Llama2. Additionally, we assess the CREAM-Chat model with Needle-in-a-Haystack [^1] https://github.com/gkamradt/LLMTest_NeedleInAHaystack and LongBench[16]. Unless otherwise specified, we use the linear interpolation to adapt LLMs to a longer context length.
@@
As far as we know, RandPos [17] and PoSE [1] are similar to our approach in that they manipulate position indices to enable fine-tuning on the pre-trained length for context expansion. Therefore, these two methods serve as the baselines for our primary comparisons. More details about the experimental setup can be found in the Appendix C.
@@ 3.2 Effective Context Window Size Evaluation on CREAM-Base
We evaluate the long-context understanding capabilities of the CREAM-Base model on two tasks: LongChat-Lines [^2] Passkey retrieval [28] is another similar task for evaluating long-context LLMs, but it is too simplistic to reflect model performance at different context window sizes, so we use the dataset provided by Pal et al. [26], which closely aligns with the task described in Li et al. [29] [26] (Figure 3) and “Lost in the Middle” [15] (LABEL:tab:lost_in_the_middle).
Figure 3: Results (%) on LongChat-Lines. Each length consists of 50 samples. All results are fine-tuned on Llama-2-7B with 4K length data through linear position interpolation. Refer to Appendix D for ablated results using NTK [8] and Yarn [9].
@@
CREAM-Base performs best in retrieving information from long context of varying lengths.
We extend the context window size up to 32K and compare CREAM with the Llama 2-7B [2], RandPos [17], and PoSE [1]. As the context window size increases, the performance of all models drops, but CREAM always performs best except for the window size of 3.6K (see Figure 3). In terms of the average performance over all context window sizes, CREAM outperforms PoSE by 16%, demonstrating its good long-context understanding ability.
@@
Lost-in-the-Middle is an observation that LLMs are generally good at retrieving relevant information appearing at the beginning/end of the input context [15]. To validate the effectiveness of our middle-focused truncated Gaussian sampling, we evaluate CREAM and compare it with PoSE on the key-value retrieval task proposed by Liu et al. [15]. We present results in LABEL:tab:lost_in_the_middle, where the cyan shading indicates middle segments. We find that: regardless of the chosen interpolation method, CREAM always outperforms PoSE by a large margin. e.g., CREAM-Linear surpasses PoSE-Linear by 21.2% when the relevant information is placed at 18.
@@ 3.3 Long Context Understanding Evaluation on CREAM-Chat
We conduct long-context evaluation of CREAM-Chat on two tasks:
  • Needle in A Haystack (Figure 9)  This task is a test that places an answer (i.e., Needle) at any position of a long context window (i.e., Haystack) and requires a model to retrieve the correct answer given a question-answer pair. We follow Wu et al. [19] and use the GPT (GPT-3.5-Turbo-0125) score as the evaluation metric.
  • LongBench (LABEL:tab:longbench2)  Bai et al. [16] is a more realistic benchmark because it covers real-world application scenarios like long-context QA and summarization. Moreover, it is specifically designed for Chat models.
@@
Figure 4: Results on Needle-in-a-Haystack. indicates the results excerpted from Wu et al. [19]. Both results are instruction-tuned on LLaMa2-7B-Chat with 4K length data. The color gradually changes from deep green to deep red, indicating the Recall performance decreases from 10 to 1.
Table 2: Results (%) on LongBench. indicates results reported by Bai et al. [16]. CREAM-7B-32k is instruction-tuned for 100 steps using 4K length data on LLaMa2-7B-Chat.
@@
We visualize the results of CREAM-Chat and the recent SkipAlign in Figure 9. Clearly, CREAM-Chat beats SkipAlign because the performance of SkipAlign [19] decreases from the window size of 18K while CREAM-Chat displays a perfect performance everywhere until from the window size of 29K. Notably, CREAM-Chat is only fine-tuned for 100 steps.
@@
We present results on LongBench in LABEL:tab:longbench2. CREAM-Chat again surpasses strong baseline models, demonstrating its better use of extended context size. In terms of the average performance over all tasks, it outperforms the second best model, i.e., LongChat-v1.5-7B-32k [29] , by 1.6%, though it is only tuned on a very small amount of data and for only 100 steps.
@@ 3.4 Language Modeling and Standard Benchmark
Following Chen et al. [7], Zhu et al. [1], Peng et al. [9], we perform the classic language modeling evaluation, i.e., perplexity evaluation, on GovReport [30] and Proof-pile [31]. Since a lower perplexity does not necessarily imply better model performance on downstream tasks [11, 32, 33, 34], we further conduct evaluation on the standard natural-language-understanding (NLU) benchmark [27]. This also lets us know whether fine-tuning hurts the NLU ability of the pre-trained base model.
@@
We apply different positional interpolation methods to RandPos [17], PoSE [1], and CREAM and report their perplexities in LABEL:tab:linear_ppl. We find that: CREAM and PoSE have a similar perplexity in different settings and both outperform RandPos. This occurs primarily because the position indices used during RandPos fine-tuning are discontinuous, which creates an inconsistency with the pre-training stage.
Table 4: Experimental results of standard benchmarks. indicates results cited from Touvron et al. [2], and all results are based on LLaMa2-7B fine-tuned on 4K data length.
@@
Ideally, fine-tuning should not adversely affect the original capabilities of the pre-trained base model. Our evaluation of CREAM confirms this, i.e., CREAM nearly retains all NLU abilities of the base Llama2-7B (see LABEL:tab:standard_benchmark). Interestingly, CREAM improves over Llama2-7B on ARC-C and HellaSwag. This is because these two tasks are few-shot tasks with longer prompts, necessitating the assistance of long-context understanding. Again, PoSE [1] performs similar to CREAM and both beats RandPos [17].
@@
We push the limit and extend the context length of Llama-2-7B up to 256K. Following Zhu et al. [1], we evaluate the extended model by calculating the average perplexity over 20 samples from PG-19 [35] and Book3 [36]. [^3] We use sliding window for calculation, with a window size of 32,768 and a sliding step size of 4,096. Since the PG-19 test set does have enough samples that are longer than 256K, we select a subset of samples from the PG-19 training set.
Table 5: Perplexity results of PG-19 and Book3. indicates results copied from Zhu et al. [1], and CREAM is based on LLaMa2-7B fine-tuned on 4K data length.
@@
We experiment with target context lengths 64K, 96K, 128K, 192K, and 256K and apply different positional interpolation methods to the extended model (see LABEL:tab:extremely_long). The results of PoSE [1] in LABEL:tab:extremely_long are based on fine-tuning LLaMa 1-7B with 2K data length, and are provided for reference only. Surprisingly, the increase of the target context length brings little to no perplexity increase, demonstrating the stability of CREAM across different target context lengths, even when the target context is extremely long.

### v2

# An Efficient Recipe for Long Context Extension via Middle-Focused Positional Encoding

@@ Abstract
Recently, many methods have been developed to extend the context length of pre-trained large language models (LLMs), but they often require fine-tuning at the target length ($\gg 4K$) and struggle to effectively utilize information from the middle part of the context. To address these issues, we propose Continuity-Relativity indExing with gAussian Middle (CREAM), which interpolates positional encodings by manipulating position indices. Apart from being simple, CREAM is training-efficient: it only requires fine-tuning at the pre-trained context window (e.g., Llama 2-4K) and can extend LLMs to a much longer target context length (e.g., 256K). To ensure that the model focuses more on the information in the middle, we introduce a truncated Gaussian to encourage sampling from the middle part of the context during fine-tuning, thus alleviating the “Lost-in-the-Middle” problem faced by long-context LLMs. Experimental results show that CREAM successfully extends LLMs to the target length for both Base and Chat versions of Llama2-7B with “Never Miss A Beat”. Our code is publicly available at https://github.com/bigai-nlco/cream.
@@
Figure 1: Results of applying different position interpolation methods to the “Lost-in-the-Middle” task on CREAM and PoSE (Zhu et al., 2023). We can see that CREAM outperforms PoSE (Zhu et al., 2023) at every position, with a particularly improvement in the middle.
@@ 1 Introduction
Transformer-based Large Language Models (LLMs) are typically pre-trained with a fixed context window size, e.g., 4K tokens in Touvron et al. (2023a). However, many downstream applications, including in-context learning (Huang et al., 2023; Li et al., 2023a) and LLM agents (Qian et al., 2023; Zheng et al., 2023) necessitate the processing of significantly longer contexts, e.g., up to 256K tokens. Recent works have proposed promising approaches that efficiently extend the context window of pre-trained LLMs by interpolating Positional Encodings (PEs) (Chen et al., 2023; Peng and Quesnelle, 2023; Peng et al., 2023; Xiong et al., 2023; Zhang et al., 2024) with a short period of fine-tuning. Unlike other techniques such as efficient transformer (Tworkowski et al., 2024; Munkhdalai et al., 2024) and memory augmentation (Tan et al., 2024), PE-based methods do not necessitate alterations to the model’s architecture or the incorporation of supplementary modules. Consequently, PE-based methods offer the advantages of straightforward implementation and rapid adaptation, making them a practical solution for extending the operational range of LLMs in tasks involving larger context windows.
Despite the simplicity and effectiveness, existing PE-based methods exhibit two significant limitations. First, prior approaches, such as positional interpolation (Chen et al., 2023), still require fine-tuning on the target context window size, which imposes a substantial computational overhead (Zhu et al., 2023). Secondly, though some PE methods have demonstrated potential in handling extremely long sequences, as evidenced by low sliding window perplexity scores, their performance deteriorates notably in “in-the-middle” scenarios (Liu et al., 2024). Specifically, when the model is required to accurately retrieve and process content located in the middle of an extended context, there is a marked drop in performance on the extended window size (Figure 1 and Figure 3).
@@
    CREAM can not only fine-tune within the pre-training context window size, but also alleviate the issue of the model easily getting lost in the middle. e.g., CREAM-YaRN outperforms PoSE-YaRN  (Zhu et al., 2023) by over 20% on average in the “Lost in the Middle” (Liu et al., 2024) task.
@@
    CREAM can further be enhanced by integrating novel designs on positional interpolation frequencies (such as Linear (Chen et al., 2023), NTK (Peng and Quesnelle, 2023), Yarn (Peng et al., 2023), etc.), and can be extended to context window sizes of up to 256K or beyond.
@@
    CREAM-Chat model requires only 100 steps of instruction-tuning to achieve nearly perfect performance on the Needle-in-a-Haystack pressure test, and it outperforms existing strong baselines on LongBench (Bai et al., 2023).
@@ 2.1 Preliminaries
Transformer-based language models typically encode positional indices sequentially as $\{0,1,\cdots,N-1\}$. Traditional length extension methods (Chen et al., 2023; Peng and Quesnelle, 2023; Peng et al., 2023) directly fine-tune on the target length $L$ with an updated positional index. This approach preserves the continuity of all absolute positions and learns all position indices within $[0,L-1]$, thereby successfully extending to the target length. Furthermore, PoSE (Zhu et al., 2023) attributed their superior performance over RandPos (Ruoss et al., 2023) to the ensured continuity of segments during fine-tuning.
@@
Relative positional encoding (RPE) (Shaw et al., 2018) has been proposed as an effective positional encoding method, where only the relative positions between two tokens are considered. Similar to prior works (Ruoss et al., 2023; Zhu et al., 2023; Wu et al., 2024), our work focuses on rotary positional encoding (RoPE) (Su et al., 2024), which is one of the most prominent RPE methods and has been widely applied to LLMs including the recent Llama family (Touvron et al., 2023b, a; AI@Meta, 2024). In RoPE, only the relative distances between position pairs $(|j-i|;0\leq i<j\leq L-1)$ are learned during fine-tuning (Appendix A). Due to this property, we can manipulate the position indices such that all relative positions between $[0,L-1]$ are learnable within the pre-trained window size.
@@ 2.2 Proposed Recipe: Continuity-Relativity indExing with gAussian Middle (CREAM)
We first discuss the motivations behind our design of the context length. First, prior works (Han et al., 2023; Xiao et al., 2023) observed that a significant amount of attention score is allocated to the beginning tokens of a sequence, which can potentially encode absolute positional information even without explicit positional encoding (Kazemnejad et al., 2024). Secondly, the starting and ending tokens of long contexts can be treated as two pointers that localize the middle indices with the help of relative encodings. Therefore, we divide the pre-trained context window into three segments. The detailed ablation results are shown in Section 3.6.
@@
For the sake of continuity, we set the $L_{h}$ and $L_{t}$ to a very small value $k$, where $0\textless k\ll N$. Specifically, we use $k=32$ in our experiments. This choice allows the middle segment to closely approximate the pre-trained context window. To maintain relativity, we divide $N$ equally into three parts and fix the $L_{h}$ and $L_{t}$ to $N/3$, enabling the model to learn as many relative positions as possible. In our fine-tuning process, both types of examples are sampled with equal probability to maintain balance.
@@
To better focus the training process on the middle part of the long context, we introduce a truncated Gaussian function. This approach reduces the interval overlap in Equation 2 and directs the model’s attention toward the middle section of the long context. In Appendix B, we provide theoretical justifications of our truncated Gaussian design, indicating that the maximization of $|D_{r}|$ holds for middle positions in $[N,L/2)\cup(L/2,L-N]$.
@@ 3.1 Experimental Setup
We use Llama-2-7B and Llama-2-7B-Chat (Touvron et al., 2023a) as the base models and extend their pre-trained context window size of 4K to a target context length of 32K. The extended models are referred to as CREAM-Base and CREAM-Chat, respectively. Note that, though the target context length is 32K, we do not have to fine-tune CREAM on 32K token long text (see Section 2.2).
@@
We conduct long-context LLM evaluation of CREAM-Base on LongChat-Lines (Pal et al., 2023) and Lost-in-the-Middle (Liu et al., 2024). Ideally, fine-tuning should not disrupt what the base model has learned, so we further evaluate CREAM-Base on the language modeling task and the evaluation benchmark (Beeching et al., 2023) adopted by Llama2. Additionally, we assess the CREAM-Chat model with Needle-in-a-Haystack [^1] https://github.com/gkamradt/LLMTest_NeedleInAHaystack and LongBench(Bai et al., 2023). Unless otherwise specified, we use linear interpolation to adapt LLMs to a longer context length.
@@
As far as we know, RandPos (Ruoss et al., 2023) and PoSE (Zhu et al., 2023) are similar to our approach in that they manipulate position indices to enable fine-tuning on the pre-trained length for context expansion. Therefore, these two methods serve as the baselines for our primary comparisons. More details about the experimental setup can be found in the Appendix C.
@@ 3.2 Effective Context Window Size Evaluation on CREAM-Base
We evaluate the long-context understanding capabilities of the CREAM-Base model on two tasks: LongChat-Lines [^2] Passkey retrieval (Mohtashami and Jaggi, 2023) is another similar task for evaluating long-context LLMs, but it is too simplistic to reflect model performance at different context window sizes, so we use the dataset provided by Pal et al. (2023), which closely aligns with the task described in Li et al. (2023b) (Pal et al., 2023) (Figure 3) and “Lost in the Middle” (Liu et al., 2024) (LABEL:tab:lost_in_the_middle).
Figure 3: Results (%) on LongChat-Lines. Each length consists of 50 samples. All results are fine-tuned on Llama-2-7B with 4K length data through linear position interpolation. Refer to Appendix E for ablated results using NTK (Peng and Quesnelle, 2023) and Yarn (Peng et al., 2023).
@@
CREAM-Base performs best in retrieving information from long contexts of varying lengths.
We extend the context window size up to 32K and compare CREAM with the Llama 2-7B (Touvron et al., 2023a), RandPos (Ruoss et al., 2023), and PoSE (Zhu et al., 2023). As the context window size increases, the performance of all models drops, but CREAM always performs best except for the window size of 3.6K (see Figure 3). In terms of the average performance over all context window sizes, CREAM outperforms PoSE by 16%, demonstrating its good long-context understanding ability.
@@
Lost-in-the-Middle is an observation that LLMs are generally good at retrieving relevant information appearing at the beginning/end of the input context (Liu et al., 2024). To validate the effectiveness of our middle-focused truncated Gaussian sampling, we evaluate CREAM and compare it with PoSE on the key-value retrieval task proposed by Liu et al. (2024). We present results in LABEL:tab:lost_in_the_middle, where the cyan shading indicates middle segments. We find that: regardless of the chosen interpolation method, CREAM always outperforms PoSE by a large margin. e.g., CREAM-Linear surpasses PoSE-Linear by 21.2% when the relevant information is placed at 18.
@@ 3.3 Long Context Understanding Evaluation on CREAM-Chat
We conduct long-context evaluations of CREAM-Chat on two tasks:
  • Needle in A Haystack (Figure 10)  This task is a test that places an answer (i.e., Needle) at any position of a long context window (i.e., Haystack) and requires a model to retrieve the correct answer given a question-answer pair. We follow Wu et al. (2024) and use the GPT (GPT-3.5-Turbo-0125) score as the evaluation metric.
  • LongBench (LABEL:tab:longbench2)  Bai et al. (2023) is a more realistic benchmark because it covers real-world application scenarios like long-context QA and summarization. Moreover, it is specifically designed for Chat models.
@@
Figure 4: Results on Needle-in-a-Haystack. indicates the results excerpted from Wu et al. (2024). Both results are instruction-tuned on LLaMa2-7B-Chat with 4K length data. The color gradually changes from deep green to deep red, indicating the Recall performance decreases from 10 to 1.
Table 2: Results (%) on LongBench. indicates results reported by Bai et al. (2023). CREAM-7B-32k is instruction-tuned for 100 steps using 4K length data on LLaMa2-7B-Chat.
@@
Mistral-7B-Instruct-v0.1 29.5 20.7 26.4 13.6 29.6 10.8 21.8
Mistral-7B-Instruct-v0.2 28.5 21.5 26.1 50.1 33.8 13.9 29.0
Mistral-7B-Instruct-v0.3 33.2 30.6 26.8 56.4 15.3 10.4 28.8
@@
We visualize the results of CREAM-Chat and the recent SkipAlign in Figure 10. Clearly, CREAM-Chat beats SkipAlign because the performance of SkipAlign (Wu et al., 2024) decreases from the window size of 18K while CREAM-Chat displays a perfect performance everywhere until from the window size of 29K. Notably, CREAM-Chat is only fine-tuned for 100 steps.
@@
We present results on LongBench in LABEL:tab:longbench2. CREAM-Chat again surpasses strong baseline models, demonstrating its better use of extended context size. In terms of the average performance over all tasks, it outperforms the second best model, i.e., LongChat-v1.5-7B-32k (Li et al., 2023b), by 1.6%, though it is only tuned on a very small amount of data and for only 100 steps.
3.4 Effectiveness of PEFT Integration
To demonstrate that CREAM can be directly combined with PEFT techniques (such as LoRA (Hu et al., 2022) and QLoRA (Dettmers et al., 2023)), requiring no additional modifications. We conducted experiments on LLaMa-2-7B-Chat using the identical dataset and settings. The experimental results are presented in Table LABEL:tab:lora. The results indicate that models fine-tuned using LoRA and QLoRA achieve performance nearly equivalent to those fine-tuned with full parameter.
Table 3: Results (%) on LongBench. indicates results reported by Bai et al. (2023). CREAM-7B-32k is instruction-tuned for 100 steps using 4K length data on LLaMa2-7B-Chat.
Model Single- Doc QA Multi- Doc QA Summari- zation Few-shot Learning Code Completion Synthetic Tasks Macro
Llama2-7B* 24.9 22.6 24.7 60.0 48.1 5.9 31.0
LoRA 28.9 28.6 27.8 62.2 54.6 10.8 35.5
QLoRA 28.1 27.6 28.1 61.7 54.6 10.1 35.0
CREAM-7B-32k 34.8 31.1 27.2 65.1 50.4 7.0 35.9
@@ 3.5 Language Modeling and Standard Benchmark
Following Chen et al. (2023); Zhu et al. (2023); Peng et al. (2023), we perform the classic language modeling evaluation, i.e., perplexity evaluation, on GovReport (Huang et al., 2021) and Proof-pile (Zhangir Azerbayev, 2022). Since a lower perplexity does not necessarily imply better model performance on downstream tasks (Zhang et al., 2024; Hu et al., 2024; Arora et al., 2024; Park et al., 2024), we further conduct evaluation on the standard natural-language-understanding (NLU) benchmark (Beeching et al., 2023). This also lets us know whether fine-tuning hurts the NLU ability of the pre-trained base model.
@@
We apply different positional interpolation methods to RandPos (Ruoss et al., 2023), PoSE (Zhu et al., 2023), and CREAM and report their perplexities in LABEL:tab:linear_ppl. We find that: CREAM and PoSE have a similar perplexity in different settings and both outperform RandPos. This occurs primarily because the position indices used during RandPos fine-tuning are discontinuous, which creates an inconsistency with the pre-training stage.
Table 5: Experimental results of standard benchmarks. indicates results cited from Touvron et al. (2023a), and all results are based on LLaMa2-7B fine-tuned on 4K data length.
@@
Ideally, fine-tuning should not adversely affect the original capabilities of the pre-trained base model. Our evaluation of CREAM confirms this, i.e., CREAM nearly retains all NLU abilities of the base Llama2-7B (see LABEL:tab:standard_benchmark). Interestingly, CREAM improves over Llama2-7B on ARC-C and HellaSwag. This is because these two tasks are few-shot tasks with longer prompts, necessitating the assistance of long-context understanding.
@@
We push the limit and extend the context length of Llama-2-7B up to 256K. Following Zhu et al. (2023), we evaluate the extended model by calculating the average perplexity over 20 samples from PG-19 (Rae et al., 2019) and Book3 (Presser, 2020). [^3] We use sliding window for calculation, with a window size of 32,768 and a sliding step size of 4,096. Since the PG-19 test set does have enough samples that are longer than 256K, we select a subset of samples from the PG-19 training set.
Table 6: Perplexity results of PG-19 and Book3. indicates results copied from Zhu et al. (2023), and CREAM is based on LLaMa2-7B fine-tuned on 4K data length.
@@
We experiment with target context lengths 64K, 96K, 128K, 192K, and 256K and apply different positional interpolation methods to the extended model (see LABEL:tab:extremely_long). The results of PoSE (Zhu et al., 2023) in LABEL:tab:extremely_long are based on fine-tuning LLaMa 1-7B with 2K data length, and are provided for reference only. Surprisingly, the increase of the target context length brings little to no perplexity increase, demonstrating the stability of CREAM across different target context lengths, even when the target context is extremely long.